a 1 1 1 1 a 1 9. Find Write the answer in completely factored form. 1 a 1 1 1 a
Answer #1

Page No. Date a 1 1 a - - - a a a 1-14 + a ++ +1 a 1 - a - -1 = a 9(+1)-1 (0-1)+1 (1-4) Tira-13-1 (0-1)+161-4) +11) (0-1) – aDate q 0 - 2 (0-1 (2-a +قه (6-1) كه] - الا-ه (تف = ه دوم (1-0) - 4 -) 0+ a- za) (1-1) - (a-.) = [ata-2a -3 (4-11]a=1) ] (2-1)Page No. Date I a = (2+3) (a-1)3 + ta 1 a